OpenStudy (blackops2luvr):

The sets of numbers 6, 8, 10 and 5, 12, 13 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples! ~iGreen

OpenStudy (blackops2luvr):

@iGreen

OpenStudy (blackops2luvr):

Last question, that's why its hard <3

OpenStudy (igreen):

We can plug them into the Pythagorean Theorem. \(a^2 + b^2 = c^2\) \(6^2 + 8^2 = 10^2\) \(36 + 64 = 100\) 36 + 64 = ? @blackops2luvr

OpenStudy (blackops2luvr):

,lol... its, soooooo hard, i dont know.... i cant....

OpenStudy (blackops2luvr):

its 100

OpenStudy (blackops2luvr):

so, i out in everything u just typed?

OpenStudy (blackops2luvr):

i put*

OpenStudy (igreen):

Yep, so that gives us 100 = 100, which means 6, 8, and 10 is a true Pythagorean triple because it satisfies the equation \(a^2 + b^2 = c^2\).

OpenStudy (blackops2luvr):

thank's

OpenStudy (igreen):

Now we plug the 2nd one into the Pythagorean theorem: \(a^2 + b^2 = c^2\) \(5^2 + 12^2 = 13^2\) \(25 + 144 = 169\) 25 + 144 = ? @blackops2luvr

OpenStudy (blackops2luvr):

169.

OpenStudy (blackops2luvr):

so now i type tht?

OpenStudy (igreen):

Yep, so that gives us 169 = 169. So 5, 12, and 13 is a perfect Pythagorean Triple too because it satisfies the equation \(a^2 + b^2 = c^2\).

OpenStudy (igreen):

Just write something like this: Let's plug in 6, 8 and 10 into the pythagorean theorem: a^2 + b^2 = c^2 6^2 + 8^2 = 10^2 36 + 64 = 100 100 = 100 So 6, 8, and 10 are pythagorean triples because they satisfy the pythagorean theorem. Let's plug in 5, 12 and 13 into the pythagorean theorem: a^2 + b^2 = c^2 5^2 + 12^2 = 13^2 25 + 144 = 169 169 = 169 So 5, 12, and 13 are also pythagorean triples because they satisfy the pythagorean theorem.

OpenStudy (blackops2luvr):

wow. thank you soooooooooooooooooooooooooooooooooooooooooo much dude,

OpenStudy (blackops2luvr):

the test is good to submit rght?

OpenStudy (igreen):

Yes, I think so..tell me what score you get.

OpenStudy (blackops2luvr):

i got a 80% ~one essay ungraded

OpenStudy (blackops2luvr):

i got everything correct.

OpenStudy (igreen):

Cool.

OpenStudy (blackops2luvr):

well... i go now. I neeeeeed to play my Advanced Warfare :D

OpenStudy (blackops2luvr):

bye!

OpenStudy (igreen):

Bye!